Find the equation of line which bisects the join of (2,-5) and (6,3) and bisects the join of (-1.1) and (-5,7)


1
Expert's answer
2021-12-07T09:48:43-0500

(a) Let us call points AB

The right bisector of a line segment bisects the line segment at 90∘



The end-points of the line segment are given as A(2,-5) and B(6,3).

"Accordingly,\ mid-point\ of\ AB = (\frac{2+6}{2},\frac{-5+3}{2})=(4,-1)"


And slope of AB"=\frac{3--5}{6-2}=2"

therefore the slope of line perpendicular to AB"=-\frac{1}{2}"


Thus equation of the line passing through (4,-1) and having a slope of "-\frac{1}{2}" is given by,

"(y+1)=-\frac{1}{2}(x-4)"


"y+\frac{1}{2}x=1"

thus the equation of the line is "y+\frac{1}{2}x=1"



(b)Let us call points CD

The right bisector of a line segment bisects the line segment at 90∘



The end-points of the line segment are given as C(-1,1) and D(-5,7).

"Accordingly,\ mid-point\ of\ CD = (\frac{-1-5}{2},\frac{1+7}{2})=(-3,4)"


And slope of CD "=\frac{7-1}{-5--1}=\frac{-3}{2}"

therefore the slope of line perpendicular to CD"=\frac{2}{3}"


Thus equation of the line passing through (-3,4) and having a slope of "\frac{2}{3}" is given by,

"(y-4)=\frac{2}{3}(x+3)"


"y-\frac{2}{3}x=6"

thus the equation of the line is "y-\frac{2}{3}x=6"